At a university, suppose 50% of students are biology majors, 25% are economics majors. Data is also collected on class: 40% are freshmen, 30% are sophomore, 20% are juniors, and the rest are seniors. Assume that there is no double majors student,

A. Consider the % of students who are neither biology nor economics majors. This % must be between _ and _.

B. If 30% of sophomores are economics majors, what % of students are sophomores and major in economics?

C. New information shows that 30% of the students are either seniors or economics majors. Of all the seniors, what % are economics majors?
Answer #1

a) Consider the % of students who are neither biology nor economics majors. This % must be between 0 % and _25%.

b) % of students are sophomores and major in economics = 0.30*0.30 = 0.09 or 9%

c) seniors = 10%

P(senior or economic) = P(senior) + P(economic) - P(both) =  0.30

0.10 + 0.25 - P(both) = 0.30

P(both) = 0.05

P(economic major | seniors) = 0.05/0.10 = 0.5